### v1

Abstract
The abundance of information on social media has reshaped public discussions, shifting attention to the mechanisms that drive online discourse. This study analyzes large-scale Twitter (now X) data from three global debates—Climate Change, COVID-19, and the Russo-Ukrainian War—to investigate the structural dynamics of engagement. Our findings reveal that discussions are not primarily shaped by specific categories of actors, such as media or activists, but by shared ideological alignment. Users consistently form polarized communities, where their ideological stance in one debate predicts their positions in others. This polarization transcends individual topics, reflecting a broader pattern of ideological divides. Furthermore, the influence of individual actors within these communities appears secondary to the reinforcing effects of selective exposure and shared narratives. Overall, our results underscore that ideological alignment, rather than actor prominence, plays a central role in structuring online discourse and shaping the spread of information in polarized environments.
keywords:
Social Media, Ideological Alignment, Polarization, Online Discourse, Online Engagement, Public Debates
Introduction
The rapid expansion of social media has fundamentally changed the way people communicate, share information, and consume content. These platforms are now central to society, serving as tools for dissemination, entertainment, and consumption of information [1, 2, 3, 4]. However, their engagement-driven business models and the dynamics of online interactions have raised concerns about their broader social impacts [5, 6], including their role in deepening polarization [7, 8, 9], propagating misinformation [10, 11, 12], and amplifying hate speech [13, 14, 15, 16, 17].
Building on these concerns, today’s digital environment is characterized by an overabundance of information, offering users an unprecedented range of choices for media consumption [18]. This shift results from the proliferation of various sources of content, fueling intense competition for user attention [19, 20, 21], amplified by the engagement-driven strategies of social media platforms. Consequently, the information ecosystem has undergone a significant transformation, leading to a fragmented audience willing to participate [22, 23] but dispersed across numerous platforms and personalized feeds [24]. Traditional media gatekeepers, once central to public discourse, no longer unilaterally set the agenda [25], diverging from previously dominant patterns of centralized media consumption [26, 27, 28, 29]. This radical transformation has led to a vast and unstructured flow of content [30] that may induce information overload [31]. Users often feel overwhelmed by the sheer volume of content when seeking information [32], complicating their ability to navigate this landscape.
In this information-saturated ecosystem, online users increasingly exhibit selective consumption behaviors, engaging primarily with content aligned with their existing beliefs [33, 34, 12] while ignoring opposing viewpoints [35, 36]. This behavior reinforces echo chambers and exacerbates polarization [37, 38, 8], in an environment where the value of news content is no longer determined solely by its intrinsic quality [39]. News consumption has become a shared social experience, with individuals exchanging links and recommendations within their networks and treating news as a form of cultural currency [40]. As a result, traditional hierarchical models of media influence struggle to explain the dynamics of public discourse in this decentralized environment [41, 42, 43], where narratives and participants compete for prominence.
Understanding these processes requires a change in perspective. Classical theories, such as agenda-setting, emphasize the transfer of issue salience from news media to the public agenda [44, 45]. However, these theories fall short of capturing the decentralized, interactive nature of digital ecosystems [46, 27], which have become the new frontier for exploring agenda-setting effects [47]. Concepts like “agendamelding” highlight how users integrate content from multiple sources to construct personalized narratives [48, 49], reflecting the decentralization of influence. Despite the democratizing potential of this bottom-up process, questions remain about what drives public discourse in this fragmented environment.
Using multiple large-scale Twitter (now X) datasets, our analysis outlines the drivers of public discourse by examining three global debates—Climate Change (COP26), COVID-19, and the Russo-Ukrainian War. We compare various actor types, such as media outlets, activists, and individual influencers, to determine whether specific categories dominate discussions and what drives online debates.
Our findings reveal a fragmented environment where no single category of actor consistently dominates discussions. Furthermore, users cluster into two polarized ideological communities, one representing the majority of users endorsing a mainstream narrative and a minority usually embracing alternative views. Notably, a user’s stance in one debate is a reliable marker of their opinion on other topics. For instance, users supporting climate action in the COP26 debate are likely to advocate for vaccination during COVID-19 and aid for Ukraine, while opponents of COVID-19 vaccination are typically climate skeptics and critics of military aid.
These results underscore that online discussions are shaped by shared ideological alignments, rather than by a top-down model dominated by traditional gatekeepers. This highlights how decentralized, bottom-up dynamics dominate public discourse, where ideological structures are pivotal in predicting the evolution of debates and the spread of specific narratives. By identifying consistent patterns in user engagement and ideological alignment, this study provides insights into the mechanisms shaping public discourse in an information-saturated landscape.
Results and Discussion
We investigate the relationship between engagement and ideological configurations across three major debates—COP26, COVID-19, and the Russo-Ukrainian War—employing data collected with specific keyword searches (see Methods). We begin by analyzing the influence of various types of actors in driving online discussions. Next, we quantify the degree of polarization surrounding different topics and examine the relationships between users’ opinions across these topics. Finally, we explore how users in different ideological communities interact with various types of actors across topics.
Decentralized Engagement: Beyond Influencer Categories
To investigate the dynamics of engagement in online discussions, we identify a set of accounts that we refer to as ‘influencers’. We first select the top 1% of accounts receiving the highest number of retweets within each debate. We further refine the selection from this subset to include only the top 50% producers of original tweets. This approach ensures that our analysis focuses on accounts that receive significant attention and contribute to the discussion by actively creating new content.
Each influencer is then manually assigned to one of six categories: Activist, International Organization/NGO, Media, Politics, Private Individual, and Other. Accounts whose tweets are unrelated to the topics of discussion are excluded from the analysis. In addition to categorizing these accounts according to their role, we annotate each account with their stance on key issues within the debates. The final sets of influencers are comprised of 561 accounts for COP26, 1577 for COVID-19, and 1166 for Ukraine. The labeling process is described more in detail in Methods and in Supplementary Section 1.1.
To understand whether the engagement that influencers receive depends on the category they belong to, we analyze the distribution of retweets received by each category per debate. This analysis enables us to assess whether certain categories of influencers are more effective in driving users’ engagement.
Figure 1: Distribution of the number of retweets received by each influencer category and for each debate. The vertical lines reported over each distribution indicate, from left to right, the 0.05, 0.5, and 0.95 quantiles. We also report the usernames of the two most retweeted accounts per debate. No category appears to dominate over the others, but rather singular prominent accounts in each debate emerge over the rest. Notable examples are @GretaThunberg in the climate debate and @KyivIndependent in the debate about the Russo-Ukrainian War.
The distributions displayed in Fig. 1 do not indicate that specific categories of actors consistently receive more retweets. Instead, all categories receive comparable attention, suggesting that the category itself may not be a determining factor for the volume of retweets. For instance, although @GretaThunberg stands out as a prominent figure in the COP26 discussion, other activists are retweeted with comparable frequency to influencers from other categories. Similarly, in the Ukraine debate, accounts in the Media category attract engagement levels similar to influencers from the other categories, with few exceptions such as @KyivIndependent—a team of independent journalists.
To statistically test these insights, we perform pairwise comparisons between the categories within each debate using Kolmogorov-Smirnov tests, which assess whether the retweet distributions for any two categories are drawn from the same population. The results, detailed in Supplementary Section 1.3, show no statistically significant differences between distributions, suggesting that the category of an influencer does not affect the number of retweets they receive. Additional results, including the prevalence of each influencer category across retweet count buckets, are provided in Supplementary Information.
To further explore the role of influencer categories, we test whether the retweet distributions align with a null model in which these labels are assigned randomly. Two randomization strategies are applied:
  1. 1.
    Reshuffling: Category labels are randomly reshuffled across influencers, to ensure each category keeps its original size.
  2. 2.
    Uniform Sampling: Categories are randomly sampled from a uniform distribution, with each having an equal probability of selection ($p=1/6$).
We compare the observed and randomized distributions using two-sided two-sample Wilcoxon rank-sum tests for both strategies. The results show no significant differences, indicating that the observed number of retweets is consistent with random assignment of categories. This reinforces the finding that an actor’s category does not appear to affect their ability to attract retweets. Detailed results, including visualizations of the randomized distributions, are provided in Supplementary Section 1.3.
Finally, we analyze users’ preferences in retweeting influencer categories. To ensure that all users in a dataset might have retweeted each category at least once, we restrict our analysis to users with at least six retweets and compute how frequently they retweet each category. We then employ the Gini Index (GI)  [50] to measure how evenly users distribute their retweets across influencer categories. The GI is a value in $[0,1]$ that quantifies inequality within a distribution: a value of 00 corresponds to minimum inequality (i.e., all categories are retweeted by the user with the same frequency), while a value of $1$ corresponds to maximum inequality (i.e., the user only retweets a single category). See Methods for its formal definition.
The resulting distributions of users’ GI, displayed in Supplementary Fig. S5, suggest that users show a preference for a limited subset of categories. For COP26, the GI distribution has a median of $0.73$ (interquartile range: $Q1=0.60,\ Q3=0.87$); for COVID-19, the median is $0.77$ ($Q1=0.65,\ Q3=0.89$); and for Ukraine, the median is $0.69$ ($Q1=0.58,\ Q3=0.80$). However, when considering the total number of retweets each category receives, we observe lower levels of inequality across all debates: $\text{GI}=0.28$ for COP26, $\text{GI}=0.59$ for COVID-19, and $\text{GI}=0.53$ for Ukraine.
The difference between the GI estimates at the user and debate level indicates that, while most users tend to favor certain categories over others, the specific subset of preferred categories varies across individuals. Users retweet different influencer categories, but these individual preferences balance out, resulting in smaller overall inequality across categories. This pattern is visible in all three debates, particularly in the case of COP26.
Ideological Persistence as a Driver of Engagement
Our analyses show that considering only influencers’ categories is insufficient to explain the drivers of user engagement on social media. As other factors such as ideological alignment [51] have been shown to influence how people interact with social media content, we now investigate the role of ideology in shaping engagement across the key debates analyzed in this study.
In line with previous research [52, 53, 51], we estimate the ideological leaning of an individual $i$ using the “latent ideology” method. This technique infers the ideological distribution within each debate by assigning each individual a scalar value $x_{i}$ representing their ideological position. These values are derived from the users’ interactions with influencers, allowing us to capture potential patterns of ideological polarization.
We begin by constructing an interaction matrix where each element corresponds to the number of times a user retweets a specific influencer. Then, the method assigns numerical scores to users so that individuals retweeting similar subsets of influencers receive similar scores. By rescaling these values to the range $[-1,+1]$, we obtain an ideological spectrum, with users at opposing ends interacting with distinct and non-overlapping sets of influencers. A detailed explanation of the technique is provided in Methods.
Following the methodology adopted in previous works [51], we focus on the top 300 most retweeted influencers from each of the three debates. We calculate the latent ideology of all users within each debate and validate the resulting scores to ensure that the distributions represent true ideological spectra (see Supplementary Section 2.1). Finally, we follow standard procedure and estimate the ideological position of each influencer as the median ideology score of their retweeters [51].
Figure 2 illustrates the latent ideology distributions for users and influencers across the three debates. To ensure consistency, we map the majority group to $-1$ and the minority group to $+1$ for all cases. Although the extent of polarization varies between debates, all distributions exhibit a clear bimodal structure, with most users clustered near the spectrum’s extremes. This observation indicates the existence of two dominant, ideologically opposing communities within each debate.
To formally evaluate the previous observation, we compute Hartigan’s dip statistic to test for unimodality (see Methods). The results are consistent across debates, with $D=0.021$ $(P<0.001)$ for COP26, $D=0.035$ $(P<0.001)$ for COVID-19, and $D=0.068$ $(P<0.001)$ for Ukraine. These values indicate a statistically significant deviation from unimodality, confirming the presence of two well-separated communities in each debate. These ideologically opposing clusters reflect the presence of two polarized communities, where users predominantly interact with influencers aligned with their ideological stance.
Figure 2: Ideological spectra of users and influencers in the COP26, COVID-19, and Ukraine debates. The majority group is mapped to $-1$, whereas the minority group to $+1$. All distributions exhibit bimodality, indicating the presence in each debate of two main communities of users who retweet distinct sets of influencers.
Having manually annotated all influencers, we can identify the ideological stances most representative of these communities. Influencers in the Majority predominantly advocate for pro-climate action, pro-vaccination, and pro-aid stances in the COP26, COVID-19, and Ukraine discussions, respectively. Conversely, influencers in the Minority are typically associated with alternative or fringe viewpoints, such as skepticism toward climate change and COVID-19 vaccination. For instance, accounts with large followings such as @JunkScience for COP26, @HighWireTalk for COVID-19, and @georgegalloway for Ukraine, were all assigned to the ideological Minority.
Consistent with previous research on online polarization [54, 55, 8, 51], our findings reveal the presence of communities supporting conflicting ideologies. This dynamic fosters user engagement with one of these opposing narratives, resulting in a recurring pattern of polarization across debates.
To evaluate the persistence of ideological segregation between different topics, we quantify the ideological overlap of users across debates. Figure 3 presents the density plot of the joint distribution $P(x,y)$ of users’ ideologies, where $P(x)$ and $P(y)$ correspond to the ideology distributions for the debates represented on the $x$ and $y$ axes, respectively. This analysis involves $N=43,392$ users with an ideology score both in the COP26 and COVID-19 debates, $N=91,647$ in COVID-19 and Ukraine, and $N=28,463$ in Ukraine and COP26.
In the figure, points along the bisector of the first and third quadrants represent users who maintain a consistent ideological stance (Majority or Minority) across both debates. Conversely, points along the bisector of the second and fourth quadrants represent users with misaligned stances between debates (e.g., Majority in one debate and Minority in the other). To account for the size imbalance between the Majority and Minority groups shown in Fig. 2, we generate each panel of Fig. 3 by first downsampling $P(x)$ and then $P(y)$, to ensure that the number of users in the Majority group matches that of the Minority group. The full distributions without downsampling are provided in Supplementary Fig. S7.
Our findings reveal a substantial overlap of ideological communities across debates. This is reflected by the high density of users in the bottom-left (Majority in both) and top-right (Minority in both) corners of the joint distributions. Interestingly, the joint distribution of Ukraine and COP26 ideologies reveals a subset of users who hold a minority, anti-aid stance while simultaneously supporting climate action, such as @ReadeAlexandra and @DeborahDupre. This divergence suggests that the strength of ideological alignment may vary depending on the topics under discussion.
Figure 3: Joint ideology spectra of users for each pair of debates, with marginal distributions on the top and right sides. The ‘Majority’ (‘Minority’) ideology is mapped to the left (right) of the $x$-axis and the bottom (top) of the $y$-axis. Colors indicate the density of users in a region, with lighter colors corresponding to a higher concentration. Most users occupy the regions on the bottom-left and the top-right of the ideology space, indicating a substantial overlap between the Majority and Minority communities across debates.
To further assess the persistence of the Majority versus Minority configuration across debates, we calculate the probability of a user maintaining the same ideological stance across different topics. Specifically, we estimate the fraction of users within the Majority or Minority group in one debate who belong to the same group in the other two debates.
Figure 4: Conditional probability of user ideology across debates. Each tile reports the fraction of users in a ‘From’ debate who belong to the same community in a ‘To’ debate. More than 90% of users with ‘Majority’ ideology on a debate (left panel) stay in the majority community of the other two. Similarly, but to a lesser extent, users with ‘Minority’ ideology on a debate (right panel) tend to stay in the minority community of the other two, especially users within the ideological minority of the COVID-19 debate.
The results, presented in Fig. 4, further support the qualitative insights derived from the joint distribution in Fig. 3. The two matrices display the fraction of users who remain in the majority (or minority) group in a debate on the $x$-axis, given that they belong to the majority (or minority) group in the debate specified on the $y$-axis. This asymmetry arises from the different fractions of majority and minority supporters in each debate. For instance, the proportion of COP26 supporters who also endorse COVID-19 vaccines may differ from the proportion of vaccine supporters who also back climate change policies. The matrices reveal that a user’s stance in one debate is highly indicative of their stance in others, with the proportion of users maintaining the same ideological leaning equaling or exceeding 90% for most debate pairs. This result is robust to using more restrictive ideology thresholds to label users as ‘Majority’ or ‘Minority’ (see Supplementary Section 2.2).
Users in the Majority group of one debate are highly likely to belong to the Majority group in another. For instance, 99% of users in the ideological Majority of the COP26 debate also belong to the Majority in the COVID-19 debate. Similar patterns are observed for users in the Minority group, albeit with slightly less consistency. Most users in the Minority of one debate tend to adopt a Minority stance in the others. Notably, the Minority group in the COVID-19 debate exhibits a high level of ideological cohesion, with approximately 90% of these users also holding a Minority position in the COP26 and Ukraine debates. In contrast, the Ukrainian debate stands out as the only case where a Minority stance does not strongly indicate corresponding positions in both other debates.
This discrepancy may be attributed to the distinct characteristics and perceived salience of the debates under study. The COVID-19 pandemic, as a global and prolonged event, likely fostered a cohesive ideological identity among users skeptical of institutional narratives. This shared identity appears to extend across debates, such as COP26, reflecting a broader distrust of mainstream scientific and institutional messaging. Conversely, the Russo-Ukrainian War, while significant, represents a geographically and politically localized event. This localized nature may result in less ideological cohesion among Minority users across debates. Additionally, the Ukraine debate’s complexity— encompassing geopolitical perspectives, humanitarian concerns, and diverse narratives—could contribute to weaker alignment with stances on other topics.
On social media, ideological alignment has a greater impact on shaping discussions than the social, political, or professional roles of individuals or organizations. No single category of actors consistently drives the discussion. Instead, our findings highlight the persistence of a Majority versus Minority ideological configuration across debates. Users tend to consistently align with one side of the spectrum, suggesting that the structure of online discourse is less about the hierarchy of different types of actors and more about the alignment of narratives with pre-existing ideological stances.
Influencer Engagement Patterns in Majority and Minority Communities
In the previous sections, we show that most influencers receive a similar number of retweets, irrespective of their category. Furthermore, users exhibit consistent behavior by retweeting influencers aligned with the Majority or Minority ideological communities across topics. Now, we aim to investigate further how the two communities differ in terms of the influencers they retweet.
To begin, we analyze whether the number of retweets an influencer receives varies depending on the ideological community they belong to. For this purpose, we employ a bootstrapping procedure to estimate the median number of retweets and retweeters attracted by influencers associated with the Majority and Minority communities. The results, presented in Fig. 5 and detailed in Supplementary Table S3, show that across all three debates, the estimates for the two groups of influencers are comparable. This indicates that individual influencers attract similar numbers of retweets and unique retweeters, regardless of their ideological alignment, and suggests that the ideological community to which an influencer belongs does not substantially impact the engagement they receive.
Figure 5: Median number of retweets and retweeters attracted by influencers of the majority and minority communities. The estimates result from a bootstrapping procedure, with error bars corresponding to the 95% bias-corrected and accelerated confidence intervals. Across all debates, the two sets of influencers receive a comparable number of retweets from a similar number of unique users.
Next, we investigate which categories are more frequently retweeted by users with a Majority or Minority stance. For this purpose, we construct a network of influencer categories for each debate and ideological stance. In these networks, two categories are connected if a user retweets influencers from both categories at least once, with the weight of each connection representing the number of distinct users who retweeted both. Figure 6 illustrates the networks generated from this procedure.
Figure 6: Networks of influencer categories for each debate. The size of a node is proportional to the number of users who retweeted that category. The width and visibility of an edge between two categories is proportional to the number of users who retweeted both. Only retweets by users in the Majority are considered on the left column, whereas on the right column, only those by users in the Minority. The resulting networks show that the users of the two communities tend to retweet different categories of influencers. This is particularly evident in the case of the COP26 and Ukraine debates, whereas differences are not as apparent in the case of COVID-19.
In general, the resulting networks indicate that users in the Majority and Minority communities tend to retweet influencers from distinct categories. This distinction is particularly clear in the COP26 debate, where $35\%$ of all retweets to influencers by users in the Minority are received by Private Individuals, while the Majority primarily engages with Activists ($21\%$) or Organizations ($28\%$). Additionally, Media accounts receive a higher relative number of retweets from users in the Minority ($35\%$) compared to those in the Majority ($23\%$), reflecting a distinction in how the two ideological communities interact with information sources.
Similar patterns emerge in the Ukraine debate, albeit less pronounced. Media and Private Individuals are the more popular influencer categories for both ideological communities, accounting for $68\%$ and $71\%$ of all retweets to influencers by Majority and Minority users, respectively. In contrast, other categories are more popular in one community than the other, such as Politics ($15\%$ of total retweets by the Majority and $5\%$ by the Minority).
Interestingly, users in the Majority and Minority of the COVID-19 debate display similar preferences, as both communities tend to retweet each category with similar frequency relative to one another. This lack of distinction indicates that no influencer category is favored more by one ideological community over the other, as we observe instead in other debates.
These findings further reinforce the conclusion that users’ engagement on social media is shaped more by ideological alignment than by influencers’ professional or social roles. This dynamic reflects the polarization observed in online discourse, where individuals selectively interact with content that aligns with their preexisting views.
Conclusions
Our study provides new insights into the interplay between user engagement and influencers’ role in major global debates on social media. By analyzing large-scale datasets from the COP26, COVID-19, and Russo-Ukrainian War debates, we demonstrate that user ideological alignment holds greater significance than an influencer’s category in shaping online discussions.
In traditional media environments, agenda-setting theories emphasize the hierarchical influence of institutional actors, such as mainstream media and political elites, in prioritizing public issues. Our findings suggest that these dynamics are significantly transformed in the context of social media. While certain users achieve prominence within specific topics, their relevance does not appear to be tied to their social or professional role, nor is it consistent across debates. Rather, the engagement they receive may arise from resonance within polarized communities.
These findings point out a shift from the top-down agenda-setting model to decentralized, bottom-up mechanisms, where users collectively amplify narratives aligned with their ideological preferences. The persistence of a Majority versus Minority ideological configuration in all debates highlights the presence of segregated and ideologically contrasting communities that extend beyond specific topics. This consistent affiliation may further limit exposure to diverse perspectives, thereby reinforcing the effects of echo chambers. Further, despite the difference in the size of their communities, the comparable engagement levels obtained by Majority and Minority influencers suggest that their popularity is not directly related to the ideological stance they assume.
While ideological alignment is a unifying factor, our analysis also reveals topic-specific variations. For example, the Ukraine debate exhibits weaker ideological cohesion compared to COP26 and COVID-19, particularly among Minority users. This difference likely reflects the varying salience and scope of these topics. The COVID-19 pandemic, as a global and prolonged crisis, appears to have fostered a cohesive ideological identity among users with alternative views, extending across debates such as climate change. In contrast, the Ukraine conflict, characterized by regionally and politically fragmented narratives, contributes to weaker ideological alignment across topics.
The decentralized nature of social media discourse fostered a new model of information spreading, challenging the traditional gatekeeping power of dominant actors. This new paradigm presents both opportunities and risks. On one hand, it democratizes the flow of information, amplifying diverse perspectives. On the other hand, it risks exacerbating polarization and misinformation, as users selectively engage with ideologically congruent content across topics.
Despite the robustness of our analyses, several limitations must be acknowledged. First, our investigation focuses on Twitter and may not fully capture user dynamics on other platforms. Additionally, the classification of users into Majority and Minority groups relies on latent ideology modeling, which, while effective, may simplify the multifaceted nature of user beliefs. Future research should extend these analyses to explore cross-platform dynamics, temporal shifts in ideological alignment, and the role of emerging fringe platforms in shaping discourse.
Methods
Data
All data used in this work were collected via the former Twitter API for academic research, using search queries specific to each dataset.
COP26
Tweets and user information related to the COP26 debate were collected using the ‘cop26’ search query, from 1 June 2021 to 14 November 2021. The dataset includes approximately 8 million English tweets by 1 million distinct users.
COVID-19
Data were collected from Twitter in the period from 1 January 2020 to 30 April 2021 using the official Twitter API through keyword search. We collected every publicly available tweet that contained one of the following keywords: vaccin*, dose*, pharma*, immun*, no-vax*, novax*, pro-vax*, provax*, antivax*, anti-vax*. The final dataset results in approximately 35 million English tweets made by 8 million unique users.
Ukraine
Tweets related to the Russo-Ukrainian conflict were gathered by tracking specific keywords and accounts from 22 February 2022 to 17 February 2023. We leveraged a publicly available dataset [56] and hydrated a random sample comprising 25% of all available tweet IDs (https://github.com/echen102/ukraine-russia). The final hydrated dataset includes approximately 85 million English tweets by 7 million distinct users.
Identifying the Influencers in a Debate
Account Selection
For each dataset, we identify a set of ‘influencers’ that shape discussions within the context of a debate. We note that, by focusing only on English tweets, the vast majority of these influential accounts will also be English-speaking.
We start by identifying the users among a dataset’s top 1% most retweeted accounts. Users whose content is extensively reshared may have a role in driving discussions on social media, as retweeting may increase the visibility of certain content and, in turn, heighten its perceived salience in the public eye. However, focusing solely on retweets may also yield accounts that have received many retweets, but only on a small number of tweets that may have gone viral. Instead, we are interested in users who have been consistently influential in the debate, by not only receiving many retweets but also producing a sufficiently large quantity of original content. For this reason, we restrict the resulting sets of influencers in each dataset to only include the top 50% producers of original tweets (i.e., tweets that are not replies, quotes, or retweets).
Before any manual removal of accounts that may have been wrongly included (e.g., if their content is unrelated to the dataset’s theme) or that could not be classified into a category (see Supplementary Section 1.1), this selection results in 603 influencers for the COP26 debate, 2168 for the COVID-19 debate, and 1385 for the Ukraine debate.
Manual Annotation
After identifying a set of influencers for each debate, we manually label each account using ad-hoc questionnaires, in a way similar to [51]. This labeling procedure enables us to associate with each user several features that are not easily retrievable or inferable from their tweets or account information.
We assign each account to one of six predefined categories based on their social, political, or professional role, namely ‘Activist’, ‘International Organization / NGO’, ‘Media’, ‘Politics’, ‘Private Individual’, and ‘Other’. These categories are shared across all debates. Additionally, we label each account with its main ideological stance in the debate, such as their support for climate action in the COP26 debate, vaccination in the COVID-19 debate, and military aid to Ukraine. We note that the influencer set for each debate is annotated independently of the others, meaning an influencer is labeled with respect to a specific debate only if they appear among the influencers of that particular debate.
The full list of questions, in addition to further details about the annotation procedure, are reported in Supplementary Section 1.1.
Latent Ideology
The latent ideology technique [52, 53] allows for estimating a synthetic opinion distribution using retweet interactions. Following a procedure detailed in prior works [51], we begin by constructing an interaction matrix $M$, where each element $M_{ij}$ is equal to the number of times user $i$ has retweeted user $j$. We exclusively focus on the interactions between users and influencers (as defined in the previous section), with matrix rows corresponding to the former and matrix columns to the latter. Additionally, we ignore all users $i$ such that $\sum_{j}M_{ij}\leq 1$. Next, we run Correspondence Analysis [57] on matrix $M$. The method requires the following steps. First, we define a new matrix $P$ obtained from the normalization of $M$ with the total number of retweets
$P=\frac{M}{\sum_{ij}M_{ij}}$
and then consider the vector of row sum $\mathbf{r}=P\mathbf{1}$ and column sum $\mathbf{c}=\mathbf{1}^{T}P$. Then, we compute the matrix of standardized residuals of $M$ as
$S=D_{r}^{-1/2}(P-\mathbf{r}\mathbf{c})D_{c}^{-1/2}$
where $D_{r}=\text{diag}(\mathbf{r})$ and $D_{c}=\text{diag}(\mathbf{c})$. Finally, we apply single value decomposition to matrix $S$ as
$S=UD_{\alpha}V^{T}$
where $UU^{T}=VV^{T}=I$ and $D_{\alpha}$ is the matrix of singular values of $S$. User ideologies are estimated as the standard row coordinates $X=D_{r}^{-1/2}U$. We focus only on the first dimension, corresponding to the largest singular value, and rescale it to $[-1,+1]$. Like previous works [53, 51], influencer ideologies are estimated as the median ideology scores of their retweeters.
Hartigan’s dip test
Hartigan’s dip test is a statistical test designed to assess the unimodality or multimodality of a distribution [58]. Specifically, it tests the null hypothesis of an unimodal empirical distribution (i.e., having a single peak) against the alternative hypothesis that it is multimodal (i.e., having more than one peak). The test computes a statistic $D$ that quantifies the extent of multimodality, with larger values corresponding to larger multimodality. The statistical significance $p$ of the test is used to assess the multimodality of the distribution, with low $p$-values indicating that the hypothesis of unimodality can be rejected. Notably, this test can be used to assess whether an empirical ideology distribution can be considered bimodal. In this work, we used this test to assess whether the user opinion distributions inferred with the latent ideology method can be considered bimodal (Fig. 2), as distributions displaying two peaks at the opposing ends of a spectrum are characteristic of a polarised debate.
Gini Index
The Gini Index (or coefficient) is a measure that quantifies the inequality in the frequency values of a distribution [50]. For a vector $\mathbf{x}$ of $n$ frequencies, it is defined as
$\text{GI}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}|x_{i}-x_{j}|}{2n^{2}\overline{x}}$
where $\overline{x}=\sum_{i=1}^{n}x_{i}/n$ is the mean frequency. The measure is defined in $[0,1]$, with $\text{GI}=0$ corresponding to perfect equality ($x_{i}=x_{j},\;\forall i\neq j$) and $\text{GI}=1$ to perfect inequality ($x_{i}=1$ and $x_{j\neq i}=0$). We use this measure to quantify the inequality of the proportion of retweets received by each category, both at the user level and the debate level.
Data availability
Twitter data are made available under the platform’s terms of service. Tweet IDs for COP26 are available on OSF (https://doi.org/10.17605/OSF.IO/NU75J). Tweet IDs for COVID-19 and Ukraine will be made available upon request.
Code availability
The R code used to compute the latent ideology scores is available on OSF (https://doi.org/10.17605/OSF.IO/NU75J). All other analyses were carried out using base R and freely available packages.
Author contribution
A.G. collected the data; E.L., A.G., N.D.M., A.B., W.Q. conceived the experiments; E.L. conducted the experiments; E.L., A.G., E.S., N.D.M., M.C., W.Q. analyzed the results; All authors wrote and reviewed the manuscript.
[…]
Acknowledgements A.G. gratefully acknowledges CY4GATE for the financial support.
Supplementary Information
1 Influencer Accounts
1.1 Manual Labeling
After selecting a set of influencers for each dataset (603 for COP26, 2168 for COVID-19, and 1385 for Ukraine) among the most retweeted and prolific tweet producers (see Methods in the main text), we manually labeled each account based on a predefined set of questions and possible answers. Each dataset’s questionnaire had the same overall structure, however we adapted some of the questions to align with the main themes of the respective datasets. To label an account, we used its current X profile, including the latest tweets, description, and self-reported location, and a random sample of original tweets authored by the account and available in the dataset.
The questionnaire, including all debate-specific questions, is as follows:
  1. 1.
    Has the account been deleted or suspended? (Yes, No)
  2. 2.
    Does the account belong to an individual or organization? (Individual, Organization)
  3. 3.
    What category does the account belong to? (Activist, International Organization/NGO, Media, Politics, Private Individual, Other)
  4. 4.
    Is the account associated with a specific country? (Yes, No)
  5. 5.
    If Yes to (4), what is the country’s name?
  6. 6.
    • COP26: Does this influencer appear in the dataset for reasons that are not directly related to climate change, climate politics, or the COP process? (Yes, No)
    • COVID-19: Does this influencer appear in the dataset for reasons that are not directly related to COVID-19 or health? (Yes, No)
    • Ukraine: Does this influencer appear in the dataset for reasons that are not directly related to the Russo-Ukrainian conflict? (Yes, No)
  7. 7.
    If No to (6)
    • COP26: Does this account have an obvious climate, or environmental focus? (Yes, No)
    • COVID-19: Does this account have an obvious health, vaccination or COVID focus? (Yes, No)
    • Ukraine: Does this account have an obvious focus on politics, geopolitics, or war? (Yes, No)
  8. 8.
    If No to (6)
    • COP26: Does this account express a general support for urgent climate action? (Yes, Neutral, No)
    • COVID-19: Does this account express a general support for COVID vaccination? (Yes, Neutral, No)
    • Ukraine: Does this account condemn the Russian invasion of Ukraine? (Yes, Neutral, No)
  9. 9.
    If No to (6)
    • COP26: Is this account critical of the COP process? (Yes, No)
    • COVID-19: Is this account critical of the vaccination process? (Yes, No)
    • Ukraine: Does this account express a general support for military aid to Ukraine? (Yes, No)
  10. 10.
    Are you certain of your classification of this account? (Yes, Not much, No)
From the final set of influencers, we exclude all accounts with answers ‘Yes’ to question 6 and ‘No’ to question 10. This results in 561 influencers for COP26, 1577 for COVID-19, and 1166 for Ukraine. Among these, we also include accounts that were suspended or deleted at the time of completing the questionnaires, provided their corresponding questions could still be answered reliably without accessing their profile and just using our datasets.
1.2 Prevalence of Influencer Categories
Figure S1: Proportion of influencers in each category. Next to each bar, the number of influencers in that category. All three debates are characterized by a similar combined fraction of influencers in the Media, Politics, Private Individual, and Other categories, while Activist and International Organization / NGO are more common in COP26.
Figure S2: Distribution of the number of original tweets posted by each influencer category and for each debate. We consider a tweet to be ‘original’ if it is not a retweet, a quote, or a reply. The vertical lines reported over each distribution indicate, from left to right, the 0.05, 0.5, and 0.95 quantiles. The distributions for influencers in the “Media” category exhibit longer tails compared to other categories across all datasets, while differences among other categories are less apparent or almost absent.
In this section, we provide additional information regarding the prevalence and activity of our set of influencer categories.
Figure S1 shows how each dataset’s influencers are distributed across categories. “Media” is the most populated category in all three datasets, particularly in the case of COVID-19. Additionally, all three debates have a similar share of influencers in the “Politics”, “Private Individual” and “Other” categories. Unlike the rest, the discussion around COP26 is characterized by a large number of influencers in the “Activist” and “International Organization / NGO” categories.
In Fig. S2, we show the distribution of original tweets (i.e., tweets that are neither retweets, nor quotes, nor replies) posted by the influencers in each category. Unlike the distribution of retweets received, shown in Fig. 1 of the main text, these distributions reveal some differences between influencer categories and across datasets. For instance, influencers in the “Media” category consistently exhibit distributions with longer tails compared to other categories. When alongside Fig. 1, this finding suggests that posting a large number of tweets does not directly translate to a larger count of received retweets. Similar considerations, though to a lesser extent, also apply to “Activists” in the COP26 and Ukraine debates, and to influencers labeled as “Other” in the COVID-19 dataset.
1.3 Engagement with Influencer Categories
Figure S3: Proportion of categories across different retweet classes. The left panels show the prevalence of each category within each retweet class. The right panels display the proportion of unique retweeters attracted by each category within each retweet class. Although certain categories are more prevalent than others, either consistently across all retweet classes or within specific ones, this prevalence closely resembles the proportion of users who retweeted them.
Figure S4: Boxplots of the observed number of retweets received by each influencer category and for each debate, compared with two different per-debate randomizations. For the ‘Shuffled’ randomization, we randomly permute the category label across all influencers; for the ‘Uniform’ randomization, we randomly sample the labels from a uniform distribution where each category has $p=1/6$ of being sampled. To compare the observed distributions with the two randomizations, we conduct a two-sample two-sided Wilcoxon rank-sum test for each pair and report the corresponding $P$-values adjusted with the Holm-Bonferroni method (see Table S2 for a full breakdown of the test results).
In the main text, we show that no influencer category consistently dominates over others and that individual influencers receive a comparable number of retweets, regardless of the category they belong to. Specifically, we describe testing this through two approaches: (i) conducting pairwise comparisons of the distributions of received retweets by all influencer categories per dataset, and (ii) comparing the observed retweet counts for each category and dataset with two randomized distributions.
Table S1 reports the results of all K-S tests carried out to compare each unique pair of influencer categories per dataset. The table includes sample sizes, the value of the test statistic $D$, and the corresponding $P$-values, which we adjust for multiple comparisons using the Holm-Bonferroni method. The results show no statistical evidence of a significant difference between any pairs of influencer categories, as indicated by the consistently large $P$-values. The sole exception is the comparison between “Politics” and “Other” in the Ukraine dataset, where $P<0.001$ and $P_{\text{adj}}=0.0071$.
To further investigate the relationship between received retweets and influencer category, we display in Fig. S3 the proportion of influencers in each category across different bins of retweet counts. In detail, following the approach of previous studies [1], we bin all influencers in each dataset into four classes based on the total number of retweets they received: “Low”, “Mid”, “High”, and “Very High”. Then, for each class, we examine the number of unique users who retweeted each category and qualitatively compare them, in proportion, with the number of influencers in each category. Our results reveal that, though certain categories are larger in size than others, either across all retweet bins or only within some, their prevalence approximately matches the proportion of users who retweeted them. In line with the results presented in the main text, this finding appears to indicate that some categories are more retweeted mainly because they are more populated, instead of their influencers consistently receiving more retweets.
For what concerns the comparison between the observed and randomized retweet counts, we display in Fig. S4 the resulting distributions and report in Table S2 the results of all statistical tests conducted to assess each comparison, including sample sizes, test statistic, and corresponding adjusted and non-adjusted $P$-values. We observe no statistical evidence of a significant difference between the true and randomized distribution, further suggesting that the number of retweets an influencer receives is not tied to the category they belong to.
Figure S5: Distribution of the Gini Index of users’ preference in retweeting influencer categories. The vertical lines indicate, from left to right, the 0.25, 0.5, and 0.75 quantiles of the distribution.
Figure S5 displays the distribution of users’ Gini Index (GI) concerning their retweeting habits. Specifically, the GI quantifies the inequality in how users distribute their retweets across different influencer categories. High GI values indicate that users favor retweeting specific influencer categories over others, whereas low values correspond to a uniform distribution of retweets across all categories. The resulting distributions suggest that users, across all debates, show a preference for a subset of categories.
2 Ideological Stance Estimation
2.1 Validation
Figure S6: Validation of the latent ideology technique. (A) Classification performance of latent ideology. Here, we categorize an influencer as “Majority” (“Minority”) if they have an ideology score below (above) the threshold, and compare these labels against the answers to the questionnaire’s Question 8, ignoring “Neutral” influencers. Performance is consistent across thresholds, highlighting the method’s robustness. The performance drop seen for COP26 and COVID-19 is due to the lack of influencers with ideological scores beyond those thresholds. (B) Top three most common countries of origin/affiliation of influencers in the Majority and Minority. The frequencies observed for the two communities are consistent with each other, indicating that the resulting ideological distribution is not being affected by geographical factors. The Ukraine dataset is an exception because of the geographical aspect of its debate.
In this section, we validate the results obtained with the ‘latent ideology’ technique. In practice, we show that the resulting ideology distributions do correspond to the ideological spectra that are characteristic of each debate. To this end, we employ the influencer annotations resulting from the manual labeling procedure detailed in Section 1.1.
In the main text, we classify an influencer as “Majority” if they have a negative ideology score and as “Minority” if positive. To validate this classification, we compare it against the answers to Question 8 of the questionnaire, ignoring ‘Neutral’ influencers. This yields an F1-score of $0.80$ for COP26, $0.86$ for COVID-19, and $0.83$ for Ukraine, demonstrating that the latent ideology is effectively capturing the ideological spectrum of each debate. Furthermore, Fig. S6A shows how performance remains consistent across different thresholds for all datasets.
Figure S6B presents the top three most common countries of origin or affiliation of Majority and Minority influencers across all datasets. The observed frequencies are approximately the same for both communities, suggesting that the ideological scores are robust to the influencers’ geographical attributes. The Ukraine dataset stands as an exception, but the observed difference is motivated by the inherently geopolitical nature of the discussion.
2.2 Majority and Minority Communities
Figure S7: Joint ideology spectra of users for each pair of debates, with marginal distributions on the top and right sides. The ‘Majority’ (‘Minority’) ideology is mapped to the left (right) of the $x$-axis and the bottom (top) of the $y$-axis. Colors indicate the density of users in a region, with lighter colors corresponding to a higher concentration. Differently to Fig. 3 in the main text, here we perform no downsampling of the Majority community and instead omit low-density regions (i.e., the panels’ white background). Nonetheless, the same ideological configuration emerges in all three panels.
Figure S8: Conditional probability of user ideology across debates. Each tile reports the fraction of users in a ‘From’ debate who belong to the same community in a ‘To’ debate. This figure extends Fig. 4 in the main text by reporting the probabilities resulting from different thresholds used to categorize users based on their ideology. In detail, we label as Majority or Minority all users with ideology scores below $-t$ and above $t$, respectively, ignoring all other users. In the main text, we use a threshold of $t=0$.
Joint ideological spectra
In the main text (Fig. 2), we show that the “Majority” and “Minority” communities are different in size, with the former comprised of a much larger number of users—thus, influencers—than the latter. Due to this size imbalance, the visualization of the joint ideological spectra $P(x,y)$ for all debate pairs in Fig. 4 of the main manuscript involves downsampling the Majority community. In detail, we first identify all users with an ideology score in both debates. Next, we downsample the Majority community of $P(x)$ to match the number of users in the Minority community and then do the same for $P(y)$. The resulting joint PDF shows a clear ideological configuration across all three debates, with most users consistently positioned within the Majority or Minority. In Fig. S7, we show that this ideological configuration is not a product of the downsampling procedure: the same pattern is visible by performing no downsampling and simply omitting low-density regions.
Conditional probabilities
Figure 4 in the main text reports the conditional probabilities of users belonging to a community in a debate given that they belong to a specific community in another debate. The resulting values come from labeling users as “Majority” or “Minority” based on their ideology score and using $t=0$ as the categorization threshold (see Section 2.1). In Fig. S8 we show that our main finding—communities persisting across debates—is robust to the specific threshold used.
Engagement
When comparing Majority and Minority influencers in terms of the number of retweets and retweeters they attract, we avoid direct comparison of their distributions due to the large difference in sample sizes. Instead, we employ a bootstrapping procedure to estimate the median of each distribution. The results, visualized in Fig. 5 of the main text and reported in full in Table S3, suggest no substantial difference between the influencers of the two communities, as indicated by the largely overlapping estimates.
3 Tables
Table S1: Two-sided K-S test between all pairs of distributions of the number of retweets received by an influencer category. The distributions are displayed in Figure 1 of the main text. For each pair, we report the sample sizes $N_{1}$ and $N_{2}$, the test statistic $D$ and the associated $P$-value, which we also adjust for multiple comparisons per dataset using the Holm-Bonferroni method. Statistically significant results at $0.05$ are underlined. Among these, only the comparison between “Other” and “Politics” in the Ukraine dataset is significant after adjusting the $P$-values.
Dataset Category 1 Category 2 $N_{1}$ $N_{2}$ $D$ $P$ $P_{\text{adj}}$
COP26 Activist Int. Org. / NGO 77 136 0.13665011 0.31759350 1.00000000
Activist Media 77 151 0.08789886 0.82562733 1.00000000
Activist Other 77 59 0.16244772 0.29594288 1.00000000
Activist Politics 77 56 0.12012987 0.68379662 1.00000000
Activist Private Individual 77 82 0.13858093 0.37903269 1.00000000
Int. Org. / NGO Politics 136 56 0.19222689 0.09198260 1.00000000
Media Int. Org. / NGO 151 136 0.13430074 0.15130144 1.00000000
Media Other 151 59 0.14446066 0.30144968 1.00000000
Media Politics 151 56 0.15302744 0.26005408 1.00000000
Media Private Individual 151 82 0.11783234 0.45178421 1.00000000
Other Int. Org. / NGO 59 136 0.05869890 0.99651561 1.00000000
Other Politics 59 56 0.21277240 0.12186471 1.00000000
Other Private Individual 59 82 0.13662671 0.48966711 1.00000000
Private Individual Int. Org. / NGO 82 136 0.13396700 0.31756445 1.00000000
Private Individual Politics 82 56 0.22212544 0.06185157 0.92777357
COVID-19 Activist Int. Org. / NGO 30 28 0.21428571 0.44747948 1.00000000
Activist Other 30 261 0.14367816 0.58086061 1.00000000
Media Activist 676 30 0.18195266 0.29756912 1.00000000
Media Int. Org. / NGO 676 28 0.10587489 0.92383436 1.00000000
Media Other 676 261 0.07610692 0.22545267 1.00000000
Media Politics 676 163 0.07263041 0.49251884 1.00000000
Other Int. Org. / NGO 261 28 0.12712096 0.75606872 1.00000000
Politics Activist 163 30 0.20552147 0.20135056 1.00000000
Politics Int. Org. / NGO 163 28 0.09969325 0.94888753 1.00000000
Politics Other 163 261 0.13736690 0.04534181 0.68012722
Private Individual Activist 419 30 0.19928401 0.21615542 1.00000000
Private Individual Int. Org. / NGO 419 28 0.09981248 0.95621873 1.00000000
Private Individual Media 419 676 0.06416023 0.23735816 1.00000000
Private Individual Other 419 261 0.07498240 0.32640129 1.00000000
Private Individual Politics 419 163 0.07728011 0.48503396 1.00000000
Ukraine Activist Int. Org. / NGO 84 37 0.15926641 0.47469743 1.00000000
Media Activist 445 84 0.05628678 0.97858149 1.00000000
Media Int. Org. / NGO 445 37 0.17266930 0.26027967 1.00000000
Other Activist 186 84 0.16551459 0.08395698 1.00000000
Other Int. Org. / NGO 186 37 0.10389422 0.85433062 1.00000000
Other Media 186 445 0.15109339 0.00501163 0.07517448
Other Politics 186 105 0.24930876 0.00047619 0.00714280
Other Private Individual 186 309 0.10241153 0.17498476 1.00000000
Politics Activist 105 84 0.15238095 0.20622578 1.00000000
Politics Int. Org. / NGO 105 37 0.28545689 0.01758764 0.26381460
Politics Media 105 445 0.14799358 0.04840110 0.72601644
Private Individual Activist 309 84 0.08610726 0.71159914 1.00000000
Private Individual Int. Org. / NGO 309 37 0.12909997 0.64046154 1.00000000
Private Individual Media 309 445 0.06748845 0.37720526 1.00000000
Private Individual Politics 309 105 0.19325012 0.00573957 0.08609356
Table S2: Two-sample two-sided Wilcoxon rank-sum test between the observed number of retweets received by an influencer category and two randomized distributions. The true and randomized distributions are displayed in Fig. S4. For each pair, we report the sample sizes $N_{1}$ and $N_{2}$, the test statistic $W$ and the associated $P$-value, which we also adjust for multiple comparisons per dataset using the Holm-Bonferroni method. Statistically significant results at $0.05$ are underlined, however none are after adjusting the $P$-values.
Dataset Category Dist. 1 Dist. 2 $N_{1}$ $N_{2}$ $W$ $P$ $P_{\text{adj}}$
COP26 Activist True Shuffled 77 77 2961 0.99279200 1.00000000
Activist True Uniform 77 92 3515 0.93458799 1.00000000
Int. Org. / NGO True Shuffled 136 136 7944 0.04448549 0.48934039
Int. Org. / NGO True Uniform 136 96 6250 0.58152710 1.00000000
Media True Shuffled 151 151 11776 0.62114462 1.00000000
Media True Uniform 151 96 7835 0.28392097 1.00000000
Other True Shuffled 59 59 1450 0.11855345 1.00000000
Other True Uniform 59 93 2510 0.37835803 1.00000000
Politics True Shuffled 56 56 1966 0.02071237 0.24854844
Politics True Uniform 56 98 3083 0.20292026 1.00000000
Private Individual True Shuffled 82 82 3691 0.27923957 1.00000000
Private Individual True Uniform 82 86 3440 0.78737427 1.00000000
COVID-19 Activist True Shuffled 30 30 397 0.44043071 1.00000000
Activist True Uniform 30 272 3602 0.29281865 1.00000000
Int. Organization / NGO True Shuffled 28 28 449 0.35026999 1.00000000
Int. Organization / NGO True Uniform 28 231 3368 0.72137903 1.00000000
Media True Shuffled 676 676 231149 0.71090164 1.00000000
Media True Uniform 676 293 98174 0.83001698 1.00000000
Other True Shuffled 261 261 30024 0.01916287 0.22995439
Other True Uniform 261 263 31785 0.14343707 1.00000000
Politics True Shuffled 163 163 14161 0.30323547 1.00000000
Politics True Uniform 163 269 24175 0.07345626 0.80801882
Private Individual True Shuffled 419 419 90303 0.47153016 1.00000000
Private Individual True Uniform 419 249 51343 0.73322830 1.00000000
Ukraine Activist True Shuffled 84 84 3681 0.62742244 1.00000000
Activist True Uniform 84 199 8447 0.88747492 1.00000000
Int. Org. / NGO True Shuffled 37 37 560 0.18182889 1.00000000
Int. Org. / NGO True Uniform 37 177 2921 0.30344682 1.00000000
Media True Shuffled 445 445 101522 0.51290296 1.00000000
Media True Uniform 445 184 41640 0.73583023 1.00000000
Other True Shuffled 186 186 16556 0.47487687 1.00000000
Other True Uniform 186 210 18419 0.32883769 1.00000000
Politics True Shuffled 105 105 6238 0.09963063 1.00000000
Politics True Uniform 105 217 13144 0.02531641 0.30379690
Private Individual True Shuffled 309 309 45736 0.36652843 1.00000000
Private Individual True Uniform 309 179 26291 0.36359119 1.00000000
Table S3: Median number of retweets and retweeters attracted by Majority and Minority influencers. All estimates, obtained with a bootstrapping procedure with $R=50,000$ replicates, are visualized in Fig. 5 of the main text. The reported confidence intervals correspond to the 95% bias-corrected and accelerated (BCa) intervals. All estimates of mean, standard error, and confidence intervals are rounded to the nearest integer.
Dataset Estimand Community $N$ $R$ Mean Std. Error $95\%$ BCa CI
COP26 Median retweets Majority 278 50,000 2787 193 $[2426,3148]$
Median retweets Minority 22 50,000 2266 741 $[1942,4271]$
Median retweeters Majority 278 50,000 1762 130 $[1563,2039]$
Median retweeters Minority 22 50,000 1618 287 $[1252,2319]$
COVID-19 Median retweets Majority 234 50,000 9603 626 $[8716,11352]$
Median retweets Minority 66 50,000 8722 729 [7340,10131[7340,10131[ 7340 , 10131
Median retweeters Majority 234 50,000 7128 410 $[6341,7805]$
Median retweeters Minority 66 50,000 5740 496 $[4643,6510]$
Ukraine Median retweets Majority 234 50,000 10028 712 $[9052,11600]$
Median retweets Minority 66 50,000 9095 782 $[7862,10737]$
Median retweeters Majority 234 50,000 6830 391 $[5928,7413]$
Median retweeters Minority 66 50,000 5794 363 $[5162,6673]$